OpenStudy (anonymous):

the sum of two trains is 719.7 one of the trains is 4.3 faster. what are the speeds

OpenStudy (anonymous):

If you call the speed of the other train \(x\), then you can set up the equation: \[x+4.3x=719.7\] So: \[5.3x=719.7\] Can you solve this now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Divide both sides by 5.3 to get: \[x=135.79 \text{(to 2 decimal places)}\] So that's the speed of one of the trains. The question tells us that the other one is 4.3 times faster, so multiply 135.79 by 4.3 to get the speed of the other train: \[583.897\]

OpenStudy (anonymous):

its +4.3 not times

OpenStudy (anonymous):

Oh, then you will have: \[x+(x+4.3)=719.7\] So \[2x+4.3=719.7\] \[2x=715.4\] Then you can divide by 2 to find \(x\), which is the speed of the slower train, then add 4.3 onto it to find the speed of the other one.